### v2

# Autonomous Robotic Bone Micro-Milling System with Automatic Calibration and 3D Surface Fitting

Abstract
Automating bone micro-milling using a robotic system presents challenges due to the uncertainties in both the external and internal features of bone tissue. For example, during mouse cranial window creation, a circular path with a radius of 2 to 4 mm needs to be milled on the mouse skull using a microdrill. The uneven surface and non-uniform thickness of the mouse skull make it difficult to fully automate this process, requiring the system to possess advanced perceptual and adaptive capabilities. In this study, we address this challenge by integrating a Microscopic Stereo Camera System (MSCS) into the robotic bone micro-milling system and proposing a novel pre-measurement pipeline for the target surface. Starting from uncalibrated cameras, the pipeline enables automatic calibration and 3D surface fitting through a convolutional neural network (CNN)-based keypoint detection. Combined with the existing feedback-based system, we develop the world’s first autonomous robotic bone micro-milling system capable of rapidly, in real-time, and accurately perceiving and adapting to surface unevenness and non-uniform thickness, thereby enabling an end-to-end autonomous cranial window creation workflow without human assistance. Validation experiments on euthanized mice demonstrate that the improved system achieves a success rate of 85.7 % and an average milling time of 2.1 minutes, showing not only significant performance improvements over the previous system but also exceptional accuracy, speed, and stability compared to human operators.
Index Terms:
Robotics and Automation in Life Sciences, Computer Vision for Medical Robotics, Visual Servoing, Computer Vision for Automation.
I Introduction
Bone micro-milling is an essential biological operation that involves using specialized milling tools to precisely remove or shape bone tissue at a microscale level [1]. This demands a high level of skill from the human operator due to the tiny scale and the need for high accuracy and safety. Automating this process with a robotic system could, in principle, make it more accurate and reproducible than by a human operator. However, this presents challenges for the perceptual and adaptive capabilities of the robotic system. The system needs to first accurately perceive the uncertainties in the characteristics of the bone tissue, including external features such as initial pose and surface flatness, as well as internal features like thickness and heterogeneity [2], and then generate milling trajectories that can adapt to these uncertainties.
For instance, in scientific experiments involving the transplantation of human organoids into mice and monitoring their growth [3], an important preparatory step is creating a circular cranial window with a radius of 2 to 4 mm on the mouse skull, whose thickness varies unevenly from 0.27 to 0.51 mm, depending on the strain, sex, and age of the mouse. To avoid damaging the fragile membrane beneath, skilled human operators typically use a handheld microdrill to mill the skull along a circular trajectory, ensuring all positions along the trajectory reach the same expected remaining thickness simultaneously under the observation of a high magnification microscope, and then remove the central bone flap with tweezers. However, automating this process using robotic system is very complex due to the surface of the mouse skull is uneven and its thickness varies. Additionally, these features dynamically change due to the potential skull deformation caused by contact with the drill bit. Thus, the system must not only possess autonomous and accurate perception capabilities for the uneven surface and non-uniform thickness of the mouse skull, but also be able to rapidly and in real time adapt to their dynamic changes during the milling process.
Figure 1: The hardware setup of the autonomous robotic bone micro-milling system, including a robotic arm holding a microdrill, a head holder for mouse to be milled, and a MSCS.
Many studies developed robotic systems for automated cranial window creation and attempted to pre-scan the skull to obtain its surface information and thickness for milling trajectory planning. Pak et al. [4] obtains surface position and thickness by contact scanning, which involves drilling at pre-set test points while detecting changes in conductivity between the drill tip and the skull. Ghanbari et al. [5] used micro-CT, while Navabi et al. [6] combined optical coherence tomography (OCT) and machine learning, to obtain the skull anatomy. Jeong et al. [7] utilized an ultrashort pulsed Ti-sapphire oscillator to generate and detect second harmonic light for the metrology of the skull. Nonetheless, the above methods lack a feedback mechanism for real-time adaptation to the dynamic changes in the skull’s characteristics, and performing pre-scanning for each specimen may lead to a long overall duration.
TABLE I: Capabilities of this study compared to existing works.
This study [4] [5] [6] [7] [8]
Direct surface measurement ×
Real-time feedback × × × ×
Short overall duration × × ×
Without human assistance × × ×
In the prior research by the present authors [8], an autonomous robotic bone micro-milling system was developed, which employed a convolutional neural network (CNN) to process input 2D images as real-time feedback to support adaptive updates of the milling trajectory. The performance of the system was evaluated by 20 eggshell milling experiments, achieving a success ratio of 80 % and an average milling time of 16.8 minutes. However, due to the lack of direct surface measurements, the initial milling trajectory can only be set to descend from the same height at all positions, resulting in inconsistent contact times with the surface across different locations. Since no feedback signal is available before contact occurs, the system is unable to adjust the trajectory effectively during this phase, leading to increased idle time, prolonged overall milling duration, and a higher likelihood of milling failures. Moreover, before autonomous milling begins, the drill bit must be manually teleoperated to a starting position where the distance between the bit and the surface is appropriate, requiring assistance from a skilled human operator.
In this study, we introduce a Microscopic Stereo Camera System (MSCS) (see Fig. 1) and propose a pre-measurement pipeline for the target surface through automatic calibration and 3D surface fitting. The pipeline does not require a time-consuming individual pre-scanning process, allowing rapid generation of an initial trajectory without human assistance. By integrating it into the previous system, the existing real-time feedback capability was retained while adding a direct measurement to enable early perception of uneven surfaces, representing the world’s first autonomous robotic bone micro-milling system that combines pre-measurement with a real-time feedback mechanism and achieving an end-to-end autonomous cranial window creation workflow, as shown in Fig. 2. The improvement effectively reduces the idle time before contact and ensures more consistent contact timing across different positions on the milling trajectory, thereby improving both efficiency and success rate of milling process. The contrast of the capabilities of this study with existing works is shown in Table I. The main contributions of this work can be summarized as follows:
  1. 1.
    A pre-measurement pipeline was proposed and validated that, with the introduction of MSCS, starts from uncalibrated cameras to reach automatic calibration and 3D surface fitting.
  2. 2.
    The pipeline was integrated with the system in the prior research, achieving an end-to-end automation workflow.
  3. 3.
    The effectiveness and performance of the entire improved system were validated through 7 trials of autonomous cranial window creation experiments.
II Hardware Setup
Figure 2: End-to-end automation workflow of the improved autonomous robotic bone micro-milling system. The parts in black (Completion Level Recognition and Trajectory Planner) are modified from the original system and will be introduced in Section IV, while the parts in red (Automatic Calibration and 3D Surface Fitting) are newly added in this work, which are the the main contribution and will be introduced in Section III. Additionally, the robot coordinate system {R} and the microscopic 3D coordinate system {Mi} are illustrated at the bottom of the figure.
The hardware setup of the system is shown in Fig. 1, utilizing a single-arm setup from the Multiarm Robotic Platform for scientific exploration [9]. The setup includes a robotic arm (CVR038, Densowave, Japan), a linear actuator and a circular rail actuator, which can be regarded as an 8-degree-of-freedom serial manipulator with joint values $q\in\mathbb{R}^{8}$. The robotic arm is equipped with a micro drill (MD1200, Braintree Scientific, USA) designed to mill the mouse skull secured in a head holder (SG-4N, NARISHIGE, Japan). To accommodate the confined workspace of the robotic system room, a modified MSCS setup was developed based on a mobile MSCS setup [10], consisting of two cameras (STC-HD853HDMI, Omron-Sentech, Japan) fitted with low-distortion macro lenses (VS-LDA75, VS Technology, Japan). The MSCS provides a reconstruction space of approximately $20\times 30\times 10\ \mathrm{mm}$ with a depth resolution of $0.10~{}\pm~{}0.02\ \mathrm{mm}$ that is sufficient for reconstructing the mouse skull [11], and a working distance of around $500\ \mathrm{mm}$ that meets the requirements for robotic movement [9].
III Proposed Pre-measurement Pipeline
The proposed pipeline aims to automatically pre-measure the target surface and generate a 3D-fitted initial trajectory $\boldsymbol{p}_{ini,i}$ that adapts to the surface unevenness for the initialization of milling trajectory. $\boldsymbol{p}_{ini,i}$ is a $n$-dimensional vector, with each element defined by
$\boldsymbol{p}_{ini,i}\triangleq\begin{bmatrix}x_{ini,i}&y_{ini,i}&z_{ini,i}\end{bmatrix}^{\text{T}}\in\mathbb{R}^{3},$ (1)
where $i=1,2,...,n$, and $n$ refers to the the number of trajectory sampling points. The pipeline contains two parts, the Automatic Calibration (Section III-A) that determines the transformation from the microscopic 3D coordinate system {Mi} to the robot coordinate system {R} (see Fig. 2), and the 3D Surface Fitting (Section III-B) that samples trajectory points on the target surface at appropriate locations and transforms them to {R}.
Figure 3: The process of the Automatic Calibration.
III-A Automatic Calibration
With the introduction of MSCS, the point clouds of the target surface can be reconstructed with respect to {Mi} using the model in the prior research by the present authors [11]. Since the milling trajectory must be planned with respect to {R}, it is essential to determine the transformation matrix from {Mi} to {R} through calibration. However, due to the limited workspace at the microscale and insufficient features on the microdrill, the way of accessing the 3D positions of end-effector with respect to {Mi} in traditional hand-eye calibration methods become impractical, such as attaching Aruco markers [12] or checkerboard patterns [13], or identifying landmarks based on the geometric features [14]. Therefore, in this section, we propose a 2D keypoint detection-based Automatic Calibration, achieving an automatic, rapid, and real-time calibration process for microdrill without human assistance, as shown in Fig. 3. The process consists of 2D Keypoint Detection to detect the drill tip 2D positions using a CNN based on the input stereo RGB images, 3D Reconstruction to obtain 3D reconstructed drill tip points, and Landmark Calibration to solve the calibration problem between {Mi} and {R} by combining the calibration setpoints in {R} with the corresponding 3D reconstructed drill tip points in {Mi}.
III-A1 2D Keypoint Detection
To reconstruct the 3D points of the drill tip, its 2D projection coordinates in both the left and right images are required. This is primarily due to its small size, as directly accessing the reconstructed point clouds would lead to a suboptimal result. Therefore, a keypoint-detection CNN is utilized to obtain the drill tip 2D positions in both images. The CNN used is a minor adjustment of U-Net [15], where the adjustments include changing the input ground truths from segmentation masks to confidence maps of drill tip localization. By this, the drill tip 2D positions $\boldsymbol{p}_{l}\triangleq\begin{bmatrix}x_{l}&y_{l}\end{bmatrix}^{\text{T}}\in\mathbb{R}^{2}$ (pixel) in the left image and $\boldsymbol{p}_{r}\triangleq\begin{bmatrix}x_{r}&y_{r}\end{bmatrix}^{\text{T}}\in\mathbb{R}^{2}$ (pixel) in the right image can be obtained by locating the local maximum from the predicted confidence map.
III-A2 3D Reconstruction
Based on the two detected drill tip 2D positions, the 3D position of the drill tip $\boldsymbol{p}_{d}^{\{\text{Mi}\}}\in\mathbb{R}^{3}$ with respect to {Mi} can be reconstructed by:
$\boldsymbol{p}_{d}^{\{\text{Mi}\}}=\boldsymbol{M}_{d}\begin{bmatrix}\boldsymbol{p}_{l}\\ \boldsymbol{p}_{r}\end{bmatrix}+\boldsymbol{b}_{d},$ (2)
where $\boldsymbol{M}_{d}\in\mathbb{R}^{3\times 4}$ is the coefficient matrix and $\boldsymbol{b}_{d}\in\mathbb{R}^{3\times 1}$ is the bias, whose values can be calculated using the streamlined orthogonal projection model of disparity-depth mapping function (detailed in [11]):
$\boldsymbol{M}_{d}=\begin{bmatrix}P_{\rho x}&0&0&0\\ P_{\rho y}&0&0&0\\ \frac{1}{h_{\rho}}&0&-\frac{1}{h_{\rho}}&0\end{bmatrix},\ \ \boldsymbol{b}_{d}=\begin{bmatrix}-c_{x}P_{\rho x}\\ -c_{y}P_{\rho y}\\ d_{e}\end{bmatrix},$ (3)
where $c_{x},c_{y}$ (pixel) are the camera’s principal point image coordinates, $P_{\rho x},P_{\rho y}$ (pixel/mm) respectively indicate the object’s pixel sizes in the $x$ and $y$ directions, $d_{e}$ (mm) is the effective working distance of the stereo microscope setup, and $h_{\rho}$ (pixel/mm) denotes the depth response to unit disparity in reality. More details on the calibration of the camera intrinsic parameters in (3) are provided in the supplementary document. By substituting (3) into (2), the 3D position of the reconstructed drill tip can be calculated.
III-A3 Landmark Calibration
Landmark Calibration aims to determine the rigid transformation from {Mi} to {R} based on the corresponding landmarks. The transformation is defined as
$\boldsymbol{T}_{\text{\{Mi\}}\rightarrow\text{\{R\}}}\triangleq\begin{bmatrix}\boldsymbol{R}&\boldsymbol{t}\\ 0&1\end{bmatrix},$ (4)
where $\boldsymbol{R}\in\mathbb{R}^{3\times 3}$ is the rotation matrix and $\boldsymbol{t}\in\mathbb{R}^{3\times 1}$ is the translation vector.
To obtain the corresponding landmarks, the controller sends a sequence of calibration setpoints to move the drill tip, represented as $\boldsymbol{p}_{i}^{\{\text{R}\}}\in\mathbb{R}^{3},i=1,\cdots,k\in\mathbb{N}$ with respect to {R}. For each setpoint, two images are captured from left and right cameras in MSCS, and its 3D position with respect to {Mi} can be reconstructed by (2). Thus, a sequence of 3D reconstructed drill tip points with respect to {Mi} can be obtained, represented as $\boldsymbol{p}_{i}^{\{\text{Mi}\}}\in\mathbb{R}^{3},i=1,\cdots,k\in\mathbb{N}$. Note that at least $k\geq 3$ non-collinear points are required to uniquely determine the transformation.
Finally, given the source landmarks $\boldsymbol{p}_{i}^{\{\text{Mi}\}}$ and the target landmarks $\boldsymbol{p}_{i}^{\{\text{R}\}}$, the rigid transformation $\boldsymbol{T}_{\text{\{Mi\}}\rightarrow\text{\{R\}}}$ in (4) can be calculated using the VTK’s classic Landmark Transform Method [^1] https://vtk.org/doc/nightly/html/classvtkLandmarkTransform.html, thereby completing the process of Automatic Calibration. The accuracy of calibration is then evaluated by the point-to-point distances after transformation (see Section V-C).
III-B 3D Surface Fitting
In this section, the goal is to generate an initial trajectory that is 3D fitted to the surface of the target, with its center point determined through the Trajectory Center Localization and trajectory points distributed on the circular path being sampled and biased through 3D-fitted Initial Trajectory Generation.
III-B1 Trajectory Center Localization
Figure 4: Schematic of bregma and lambda, which is expected to be detected through the CNN. The milling center can be calculated by linearly weighting the two points.
In the actual experiment of mice cranial window creation, the center of the circular milling trajectory is always located between two specific anatomical points on the skull: bregma and lambda, as illustrated in Fig. 4. Inspired by the method in [16], which uses a CNN to automatically detect bregma and lambda, we applied a U-Net-based CNN with the same architecture as the one in Section III-A1 for the detection of these two points.
In practice, the ratio $l$ of the distance between the trajectory center and bregma to the distance between bregma and lambda is determined based on the radius of the milling trajectory $r$ and the size of the mouse. Therefore, the 2D position of the trajectory center in the input image can be obtained by proportionally weighting the detected positions of bregma and lambda. Then, due to the high-quality reconstruction of the skull surface, the 3D localization of the trajectory center $\boldsymbol{p}_{c}^{\{\text{Mi}\}}\triangleq\begin{bmatrix}x_{c}&y_{c}&z_{c}\end{bmatrix}^{\text{T}}\in\mathbb{R}^{3}$ with respect to {Mi} can be performed by accessing its 2D position in the RGB image from either camera and the reconstructed point cloud of the target surface (detailed in [11]).
III-B2 3D-fitted Initial Trajectory Generation
Considering the expected trajectory is circular when projected onto the $x\text{--}y$ plane, by combining the obtained trajectory center $\boldsymbol{p}_{c}^{\{\text{Mi}\}}$ and the radius of the milling trajectory $r$, the $x$ and $y$ coordinates of $n$ sampling points along the trajectory $\boldsymbol{p}_{fit,i}^{\{\text{Mi}\}}\triangleq\begin{bmatrix}x_{fit,i}&y_{fit,i}&z_{fit,i}\end{bmatrix}^{\text{T}}\in\mathbb{R}^{3},i=1,\cdots,n\in\mathbb{N}$, with respect to {Mi}, can be calculated by:
$x_{fit,i}$ $=r\cos{\theta_{i}}+x_{c}$ (5)
$y_{fit,i}$ $=r\sin{\theta_{i}}+y_{c},$
where $\theta_{i}=2{\pi}i/n,i=1,\cdots,n\in\mathbb{N}$. Then $z_{fit,i}$ can be obtained by combining $x_{fit,i},\ y_{fit,i}$ and the reconstructed point cloud of the target surface.
In order to avoid contact with the surface even before milling, an offset should be set for all sampling points along their normal directions, considering the radius of the drill tip $r_{d}$ and the potential error of Automatic Calibration. The normal vectors of all sampling points $\boldsymbol{n}_{i}\in\mathbb{R}^{3},i=1,\cdots,n\in\mathbb{N}$ can also be obtained by the reconstructed point cloud of the target surface. As a result, by substituting the result of (5), the 3D-fitted initial trajectory $\boldsymbol{p}_{ini}^{\{\text{Mi}\}}$ with respect to {Mi} can be obtained, whose each element can be calculated by
$\boldsymbol{p}_{ini,i}^{\{\text{Mi}\}}=\boldsymbol{p}_{fit,i}^{\{\text{Mi}\}}+(r_{d}+\Delta d)\cdot\boldsymbol{n}_{i}.$ (6)
where $\Delta d$ is a manually set offset for safety concerns, considering errors in Automatic Calibration and 3D Surface Fitting.
Lastly, utilizing the transformation matrix output from Automatic Calibration by (4), we can acquire the $n$-dimensional vector of transformed 3D-fitted initial trajectory $\boldsymbol{p}_{ini}$ with respect to {R}, where each element is defined by (1) and can be calculated by
$\begin{bmatrix}\boldsymbol{p}_{ini,i}\\ 1\end{bmatrix}=\boldsymbol{T}_{\text{\{Mi\}}\rightarrow\text{\{R\}}}{\cdot}\begin{bmatrix}\boldsymbol{p}_{ini,i}^{\{\text{Mi}\}}\\ 1\end{bmatrix}.$ (7)
The result of (7) is output to the Trajectory Planner (Section IV-B) for the initialization of the milling trajectory.
IV End-to-end Automation Workflow
With the integration of the proposed pre-measurement pipeline (Section III), we improve the autonomous robotic bone micro-milling system and achieve an end-to-end automation workflow (see Fig. 2). In this section, we briefly introduce the other modules of the workflow, namely Completion Level Recognition (Section IV-A) and Trajectory Planner (Section IV-B). These modules are modified from the existing system in the authors’ prior research [8] and play an essential role in providing real-time feedback after the pre-measurement phase is completed and milling begins.
IV-A Completion Level Recognition
The Completion Level Recognition aims to monitor the milling completion level along the trajectory in real-time using a CNN based on the input RGB images. The CNN follows the same DSSD-inspired [17] architecture as the previous system [8], but is retrained on a new dataset to adapt to the mouse skull milling experiment. The CNN output a bounding box for milling area detection and a heatmap image for pixel-wise completion level prediction. The network was trained on an augmented dataset derived from 587 manually collected and annotated images (a total of 18,784 training images) from videos of teleoperated robotic cranial window creation experiments conducted on 6 euthanized mice, achieving 77.6 mAP for the detection task and 24.32% MAPE for the completion prediction task, with a processing speed of 72 Hz. More network details are available in the supplementary document.
Through post-processing and sampling of the output image, an $n$-dimensional vector $\boldsymbol{c}$ can be obtained, in which element denotes the penetration percentage at the $i$-th sampling point along the real-time-updated trajectory, defined by
$\mathbb{R}\ni c_{i}\triangleq c_{i}\left(t\right)\in\left[0,1\right],$ (8)
where $i=1,2,...,n$. Note that $c_{i}=0$ means untouched and $c_{i}=1$ means completely penetrated. $\boldsymbol{c}$ is output to the Trajectory Planner for real-time trajectory updates.
IV-B Trajectory Planner
Figure 5: Schemas of (a) the Trajectory Planner in the previous system, and (b) the Trajectory Planner in this study.
The Trajectory Planner is responsible for initializing, updating and smoothing the milling trajectory. The updating process remains consistent with the previous system, employing a multimodal velocity damper to control the descent speed of the drill bit at the $i$-th sampling point to proportionally reduce based on its completion level $c_{i}$, which is real-time output from the Completion Level Recognition and defined by (8).
For the initialization process, in the previous system, due to the absence of direct surface measurements, the $z$-coordinates of all sampling points are initialized to 0 (see Fig. 5-(a)). In this study, with the integration of MSCS, a 3D-fitted initial trajectory $\boldsymbol{p}_{ini,i}$ is available, which is output from the 3D Surface Fitting (Section III-B) and defined by (1). The trajectory features non-zero $z$-axis coordinate at each sampling point to adapt to the uneven surface for (see Fig. 5-(b)), enabling the drill to contact the target surface rapidly and spatially consistently.
The smoothing process also remains the same as in the previous system, where a constrained cubic spline interpolation method [18] is used to generate a continuous trajectory based on the $n$ discrete sampling points, to match the high-precision encoders used in the robotic system [9]. Additionally, we hypothesize that if all $n$ sampling points are completely penetrated, the entire smoothed trajectory will be sufficiently milled. Therefore, when $c_{i}=1\ \ \forall i=1,\cdots,n\in\mathbb{N}$, a stop signal will be triggered and sent to the controller to automatically halt the robot’s movement and the drill’s rotation, marking the completion of the milling process. More details of the Trajectory Planner are available in the supplementary document.
V Experiments
In this section, experiments are first conducted to validate the effectiveness of the Automatic Calibration and the 3D Surface Fitting, which together constitute the proposed pre-measurement pipeline. Subsequently, further experiments are conducted to evaluate the performance of the entire autonomous robotic bone micro-milling system.
V-A System configuration
The software implementation of the experiments was carried out on an Ubuntu 20.04. The robotic arm was controlled following the approach in [9]. ROS (Robot Operating System) Noetic Ninjemys was utilized for interprocess communication, while simulations were conducted using CoppeliaSim (Coppelia Robotics, Switzerland). Communication with the robot was managed via the SmartArmStack [^2] https://github.com/SmartArmStack. Dual quaternion algebra and robot kinematics were implemented using DQ Robotics [19] in Python 3.
V-B Parameter selection
To implement the proposed method for validations and experiments, the parameters mentioned in Section III and Section IV are selected based on camera calibration, actual measurements, safety considerations, and ease of calculation: the number of trajectory sampling points $n=32$, the intrinsic parameters of cameras $c_{x}=480\ \mathrm{pixel}$, $c_{y}=270\ \mathrm{pixel}$, $P_{\rho x}=0.03947\ \mathrm{pixel/mm}$ , $P_{\rho y}=0.03894\ \mathrm{pixel/mm}$, $d_{e}=0$, and $h_{\rho}=6.4\ \mathrm{pixel/mm}$, the ratio of the distance between the trajectory center and bregma to the distance between bregma and lambda $l=1/3$, the radius of the milling trajectory $r=2\ \mathrm{mm}$, the radius of the drill tip $r_{d}=0.5\ \mathrm{mm}$, the manually set offset $\Delta d=0.5\ \mathrm{mm}$, the number of calibration setpoints $k=5$ and their coordinates $\boldsymbol{p}_{1}^{\{\text{R}\}}=\begin{bmatrix}0&-20&0\end{bmatrix}^{\text{T}},\ \boldsymbol{p}_{2}^{\{\text{R}\}}=\begin{bmatrix}10&-20&0\end{bmatrix}^{\text{T}},\ \boldsymbol{p}_{3}^{\{\text{R}\}}=\begin{bmatrix}5&-10&0\end{bmatrix}^{\text{T}},\ \boldsymbol{p}_{4}^{\{\text{R}\}}=\begin{bmatrix}10&0&0\end{bmatrix}^{\text{T}},\ \boldsymbol{p}_{5}^{\{\text{R}\}}=\begin{bmatrix}0&0&0\end{bmatrix}^{\text{T}}\ \mathrm{mm}$. Additionally, the resolution of the images captured by the MSCS is set to $540\times 960$ to meet the requirement for Automatic Calibration, 3D Surface Fitting, and Completion Level Recognition.
V-C Experiment 1: Validation of Automatic Calibration
Figure 6: Examples of the input image, ground truth, and detection result for the keypoint detection CNNs on (a) the drill tip and (b) bregma and lambda. The ground truth and detection result images are confidence maps, where pixels with brighter colors indicate a higher probability of being keypoint locations (while: drill tip, red: bregma, green: lambda).
As described in Section III-A, the accuracy of Automatic Calibration is affected by the accuracy of 2D Keypoint Detection, 3D Reconstruction, and Landmark Calibration. Considering the 3D Reconstruction depends on the MSCS, whose accuracy has been evaluated in the prior research by the present authors [11], in Experiment 1, we focus on the evaluation of 2D Keypoint Detection and Landmark Calibration.
V-C1 Training of 2D Keypoint Detection model
To prepare learning data, a total of 3,294 images satisfying the requirement that the drill tip is within the field of view were captured from videos of teleoperated robotic cranial window creation experiments conducted on 10 euthanized mice. The mice were sacrificed for another research and reused in this study, so no additional ethical approval was required. For each image, a confidence map is generated by applying a 2D Gaussian distribution to the manually annotated drill tip 2D position, which represents the probability of the drill tip’s existence and serves as the ground truth (an example is shown in Fig. 6-(a)). Data augmentation techniques, including blur and noise augmentation, color adjustment, geometric transformation, and resolution alteration, resulted in a total of 16,394 images and corresponding ground truths, with 14,754 images used for training, 820 for validation, and 820 for testing.
The models were trained in a total of 30 epochs using the Adam optimizer, with the training loss computed based on the Mean Square Error (MSE) criterion. The training rate was initialized as $1\times 10^{-3}$, followed by a reduction to $1\times 10^{-4}$ in epochs 11 to 20, and further decreased to $1\times 10^{-5}$ for epochs 21 to 30. Moreover, a batch size of 4 was employed during the training phase. All models were implemented using PyTorch 1.8.0 with CUDA 11.1, based on Python 3.8, and executed on an Ubuntu 20.04 system equipped with an NVIDIA Quadro P6000 graphics card.
V-C2 Joint evaluation metrics
Firstly, we send the robot $m=20$ validation setpoints $\boldsymbol{p}_{val,i}^{\{\text{R}\}}\in\mathbb{R}^{3},i=1,\cdots,m$ for the drill tip to reach these points. MSCS captures $m$ images each from left and right camera, outputting $m$ drill tip 2D positions each $\boldsymbol{p}_{l,i},\ \boldsymbol{p}_{r,i}\in\mathbb{R}^{2},i=1,\cdots,m$ via 2D Keypoint Detection (see Section III-A1). Then we manually annotate the left and right images, obtaining $m$ ground truths each $\boldsymbol{p}_{l,i}^{(gt)},\ \boldsymbol{p}_{r,i}^{(gt)}\in\mathbb{R}^{2},i=1,\cdots,m$. Through 3D Reconstruction (see Section III-A2), $m$ points of 3D reconstructed drill tip with respect to {Mi} can be obtained. Then, by transforming these points back to {R} using the transformation calculated from Landmark Calibration (see Section III-A3), we can obtain $\boldsymbol{p}_{trans,i}^{\{\text{R}\}}\in\mathbb{R}^{3},i=1,\cdots,m$. The accuracies of 2D Keypoint Detection and Landmark Calibration are respectively evaluated by the Root Mean Square Error (RMSE) between $\boldsymbol{p}_{l,i},\ \boldsymbol{p}_{r,i}$ and $\boldsymbol{p}_{l,i}^{(gt)},\ \boldsymbol{p}_{r,i}^{(gt)}$, and between $\boldsymbol{p}_{trans,i}^{\{\text{R}\}}$ and $\boldsymbol{p}_{val,i}^{\{\text{R}\}}$.
Given the image resolution of $540\times 960$, the error in 2D Keypoint Detection is required to not exceed 10 pixels (0.04 mm). Considering the radius of the drill tip ($r_{d}=0.5\ \mathrm{mm}$) and the manually set offset ($\Delta d=0.5\ \mathrm{mm}$) in (6), the errors of Landmark Calibration is required to not exceed 1 mm.
V-C3 Result and discussion
Figure 7: Evaluation results of (a) 2D Keypoint Detection and (b) Landmark Calibration on 20 validation setpoints. The black dots, red cross and green circle or sphere respectively represent (a) the centralized $\boldsymbol{p}_{l,i},\ \boldsymbol{p}_{r,i}$, the centralized $\boldsymbol{p}_{l,i}^{(gt)},\ \boldsymbol{p}_{r,i}^{(gt)}$, the accuracy requirement of 10 pixels; (b) the centralized $\boldsymbol{p}_{trans,i}^{\{\text{R}\}}$, the centralized $\boldsymbol{p}_{val,i}^{\{\text{R}\}}$, and the accuracy requirement of 1 mm.
As shown in Fig. 7, we randomly generated 20 coordinates as validation setpoints within the reconstruction space of MSCS, and observed an RMSE of 3.18 pixels in 2D Keypoint Detection and 0.77 mm in Landmark Calibration, which meet our accuracy requirement, indicating a high overall detection and calibration precision. Specifically, as shown in Fig. 7-(a), all 20 validation points exhibited 2D Keypoint Detection errors within the threshold, indicating that the detection model demonstrated good accuracy and robustness. However, as shown in Fig. 7-(b), 4 points exceeded the error threshold in Landmark Calibration, which is likely due to the instability of the MSCS reconstruction accuracy in certain spatial regions, such as areas with sparse textures, suboptimal viewpoints, or significant depth variations. Although these localized errors are still acceptable in Experiment 1, they may lead to milling failures in the subsequent Experiment 3. To address this issue, future work will focus on improving 3D reconstruction accuracy by employing higher-resolution cameras and more robust models, thereby enhancing spatial localization performance.
V-D Experiment 2: Validation of 3D Surface Fitting
As described in Section III-B, the accuracy of 3D Surface Fitting is affected by the accuracy of Trajectory Center Localization and 3D-fitted Initial Trajectory Generation. As the performance of 3D-fitted Initial Trajectory Generation is based on the reconstruction of the target surface, which has been evaluated in [11], and the transformation obtained via Landmark Calibration, which has been evaluated in Section V-C, in Experiment 2, we focus on the evaluation of Trajectory Center Localization.
V-D1 Training of CNN model
To prepare learning data, a dataset consisting of 523 images that satisfy the requirement that the mouse skulls in the images remain unmilled was captured from videos of the same 10 euthanized mice mentioned in Section V-C, as well as 82 images from a publicly available dataset created by [16]. The creation of ground truths and data augmentation are the same as Section V-C, resulting in a total of 8,774 images, with 7,019 for training, 877 for validation, and 878 for testing (an example is shown in Fig. 6-(b)). The training condition and parameters are also as same as Section V-C.
V-D2 Metrics
Considering that the trajectory center is calculated by linearly weighting the detected bregma and lambda, with the detection errors of these two points considered independent, the error of Trajectory Center Localization can also be calculated by linearly weighting the detection RMSE of bregma and lambda. Same as Section V-C, we require the error not to exceed 10 pixels.
V-D3 Result and discussion
In the test dataset, we observed an RMSE of 1.19 pixels for bregma detection and 1.73 pixels for lambda detection. The RMSE for lambda was higher than that for bregma, possibly because bregma is more distinguishable than lambda, even for humans. By applying linear weighting to integrate the information from both, the RMSE of the trajectory center was 1.37 pixels, meeting the accuracy requirement.
V-E Experiment 3: Validation of the autonomous robotic bone micro-milling system
Figure 8: The automation flow of an experimental case. (a) The result of Automatic Calibration (Section III-A). (b) The result of 3D Surface Fitting (Section III-B). (c) The input and output of the Completion Level Recognition (Section IV-A). (d) The result after the milling process was completed and after the bone flap was manually removed.
Following the end-to-end automation workflow shown in Fig. 2, with its hardware setup shown in Fig. 1 and described in Section II, autonomous cranial window creation experiments were conducted on euthanized mice to validate the feasibility of the proposed method.
V-E1 Animal model
All experiments were performed on mice euthanized with carbon dioxide, approved by the Animal Care and Use Committee, Graduate School of Medicine, The University of Tokyo (Approval No.: A2023M042-04). The experiments were conducted in compliance with the Fundamental Guidelines for Proper Conduct of Animal Experiment and Related Activities in Academic Research Institutions (2006), MEXT Japan, and Guidelines for Proper Conduct of Animal Experiments (2006), Science Council of Japan. 7 female NOD.CB17-Prkdcscid/J mice, aged 6 to 8 weeks and weighing 20-25 g, are used in the experiments.
V-E2 Metrics
Milling success and time are used to evaluate the performance. It should be noted that milling success is defined as the bone flap (see Fig. 8-(d)) within the milling trajectory being removable manually with tweezers after the milling is completed, without damaging the membrane beneath the skull [8].
V-E3 Result and discussion
7 experiments were conducted, and 6 were successful, resulting in a success ratio of 85.7 %. The average required time for successful milling was $2.1\pm 1.8$ minutes. Compared with the 80 % of success ratio and the $16.8\pm 2.5$ minutes of milling time using the previous system [8] evaluated on eggshell model, the feasibility and efficiency of the improved autonomous robotic bone micro-milling scheme were initially proved on euthanized mice. An example of the entire flow of a successful milling case is shown in Fig. 8. The improvement in success ratio and the reduction in milling time aligned with our expectations, as the introduction of MSCS enhanced the system’s perceptual capability for uneven surfaces, avoiding the long milling time and potential failures caused by the complete reliance on real-time feedback in the previous system.
Furthermore, compared with the results of teleoperated robotic cranial window creation mentioned in Section V-C ($13.7\pm 6.6$ minutes, 3 skilled and 4 novice human operators, 10 euthanized mice) and manual cranial window creation [6] ($8.0\pm 2.1$ minutes, 2 skilled human operators, 6 anesthetized live mice), the system is demonstrated with excellent speed, repeatability, and consistency.
The error in Completion Level Recognition is considered the main cause of the failure case, indicating that the recognition accuracy of the CNN relying solely on RGB images limits the upper bound of the success ratio. Future work will require improving the recognition accuracy of the CNN or adding additional sensors to assist with image processing.
VI Discussion and Conclusions
In this paper, we proposed a pre-measurement pipeline of automatic calibration and 3D surface fitting, and integrate it into an autonomous robotic bone micro-milling system, enabling an end-to-end autonomous cranial window creation workflow capable of rapidly, in real-time, and accurately perceiving and adapting to the uneven surface and non-uniform thickness of bone tissue without human assistance. In the experiments, we validated the system’s excellent performance and efficiency by successfully performing autonomous cranial window creation on euthanized mice, achieving faster, more accurate and more stable results compared to skilled human operators.
Future works include further improving the performance of the system and applying the system on anesthetized live mice. In addition, considering the similar challenges of target surface and thickness uncertainties in many surgeries, such as endoscopic endonasal surgery and neural surgery, we also hope to apply our autonomous robotic bone micro-milling system to other medical fields.
Acknowledgment
We would like to express our special gratitude to Dr. Saúl Alexis Heredia Pérez, Dr. Murilo Marques Marinho and the surgeons in Institute of Science Tokyo for their precious advice and kind help.